### v1

@@ 5.2 Results
Quantitative Results. To benchmark our model, we compare it against two recent methods: the unsupervised learning model CA-SUM [1] and the supervised learning model CSTA [22]. For both CA-SUM and CSTA, we follow the original implementation, downsampling each video to 2fps and using the pool5 layer of GoogleNet to extract frame-level features. Then, we use the pre-trained weights on TVSum and SumMe to predict the frame importance scores and select shots to make 15-second ads accordingly. We report the average performance on all videos.
@@
Ablation Item AP AUROC $\sigma$ $\tau$
Loss Function
CSTA [22] 0.773 0.646 0.241 0.199
@@
Window Size
CSTA (3HWS) + BCE 0.781 0.652 0.251 0.207
@@
AdSum (early fusion) 0.783 0.665 0.273 0.224
@@
Table 3 documents the results of our ablation studies. First, replacing the mean squared error loss used in [22] with binary cross-entropy (BCE) loss leads to improved performance in terms of AP, AUROC, and $\sigma$. Second, incorporating features from surrounding frames further enhances model performance. Notably, using a half window size of 3 consistently outperforms a window size of 5 across all four metrics. These findings underscore the effectiveness of our proposed 3D-CNN clip-level feature embedding in capturing temporal information.
@@ 6 Conclusion
Automated video ad clipping from long video ads into shorter versions is crucial for advertisers aiming to adapt content effectively across diverse formats while minimizing costs. To address this need, we introduce a novel ad clipping task and develop a dedicated dataset AdSum204, which consists of 204 video ads with temporally synchronized visual and audio information. Our proposed two-stream audio-visual fusion model demonstrates the effectiveness of integrating visual and audio modalities for predicting the importance of video frames. Our findings highlight the superiority of 3D CNNs for feature embedding, surpassing traditional 2D CNN-based approaches [e.g., 22]. Moreover, the early fusion strategy, which integrates both visual and audio cues, achieves the best performance. Furthermore, early fusion strategies that combine both visual and audio cues deliver the best performance, emphasizing the importance of audio in video ad clipping.

### v2

@@ 5.2 Results
Quantitative Results. To benchmark our model, we compare it against two recent methods: the unsupervised learning model CA-SUM [apostolidis2022summarizing] and the supervised learning model CSTA [son2024csta]. For both CA-SUM and CSTA, we follow the original implementation, downsampling each video to 2fps and using the pool5 layer of GoogleNet to extract frame-level features. Then, we use the pre-trained weights on TVSum and SumMe to predict the frame importance scores and select shots to make 15-second ads accordingly. We report the average performance on all videos.
@@
Ablation Item Model AP AUROC $\sigma$ $\tau$
Loss Function CSTA [son2024csta] 0.773 0.646 0.241 0.199
@@
Window Size CSTA (3HWS) + BCE 0.781 0.652 0.251 0.207
@@
Fusion AdSum (early fusion) 0.783 0.665 0.273 0.224
@@
Table 3 documents the results of our ablation studies. First, replacing the mean squared error loss used in [son2024csta] with binary cross-entropy (BCE) loss leads to improved performance in terms of AP, AUROC, and $\sigma$. Second, incorporating features from surrounding frames further enhances model performance. Notably, using a half window size of 3 consistently outperforms a window size of 5 across all four metrics. These findings underscore the effectiveness of our proposed 3D-CNN clip-level feature embedding in capturing temporal information.
@@ 6 Conclusion
Automated video ad clipping from long video ads into shorter versions is crucial for advertisers aiming to adapt content effectively across diverse formats while minimizing costs. To address this need, we introduce a novel ad clipping task and develop a dedicated dataset AdSum204, which consists of 204 video ads with temporally synchronized visual and audio information. Our proposed two-stream audio-visual fusion model demonstrates the effectiveness of integrating visual and audio modalities for predicting the importance of video frames. Our findings highlight the superiority of 3D CNNs for feature embedding, surpassing traditional 2D CNN-based approaches [e.g., son2024csta]. Moreover, the early fusion strategy, which integrates both visual and audio cues, achieves the best performance. Furthermore, early fusion strategies that combine both visual and audio cues deliver the best performance, emphasizing the importance of audio in video ad clipping.